Accelerator Update for April 6, 2009 - April 8, 2009 |
Monday April 6 The day shift began with the Tevatron (TeV) cooling down on the backup cold box #1 due to a turbine failure in cold box #2, and with the Booster, Main Injector (MI), Antiproton Source (Pbar), and MI-12B in access. Operations prepared MI-65 for access at 8:36 AM. At 9:09 AM, the I- Source vacuum pump #2 tripped off and would not stay reset. Operators contacted a Proton Source technician and Preaccelerator technicians; they brought the pump back online. Operations prepared Linac for access at 9:19 AM. Water group techs reported at 9:26 AM that they had made up water for the MI LCW system and were moving on to the TeV. FESS personnel reported at 9:26 AM that they were working on the MINOS sump pump (SMP1). The sump will be off all day and night. An Electrical Engineering (EE) Support tech reported at 9:39 AM that he had placed the TeV kickers into conditioning mode. At 10:35 AM, Operations reported water dripping from a bus cover at MI-10. A water tech investigated and said the water was coming from the roof. FESS personnel will investigate and repair. Booster came out of access at 12:16 PM. Linac came out of access at 12:19 PM. Operators reported at 2:44 PM that they had search and secured MI. Operations reported at 4:08 PM that they had searched and secured Pbar transport. Operations reported status on the following work at 5:10 PM:
Operations reported at 5:34 PM that Pbar was on and in standby. At 6:09 PM, Operations reported that a Pbar kicker (PBKICK) PC was down for the night and the sweeper magnet would not be available. At 6:52 PM, a Pbar Debuncher quadrupole power supply (D:Q701) tripped off. Operators investigated and found its 480-volt breaker tripped. Operations called in a power supply tech to check the supply. The NuMI Spokesperson asked Operations at 6:58 PM to turn on the NuMI horn. Operations reported at 7:17 PM that they had established beam to Booster, MI, and the NuMI and MiniBooNE experiments. At 8:21 PM, Operations reported that they were unable to change the settings for I:MBOFF. They believed that MECAR needed a reboot. Operators contacted an EE Support technician who agreed to the reboot. It didn’t work. The tech continued to investigate. At 8:22 PM, a Mechanical Support technician reported that the D0 service elevator was not working. Power supply techs working on D:Q701 reported at 8:53 PM that they found a loose phase cable and a failed contactor, which they repaired. A Recycler Physicist reported at 9:42 PM that he was conducting a study with high intensity protons. The EE Support tech reported at 10:25 PM that he found and fixed a bug in the MECAR software. A cryo system tech at CHL requested at 10:37 PM that operators check the TeV sector F1 wet engine. The operators reported seeing belt residue and a small wobble at one of the bearings. The engine was shut off for maintenance. ![]() A power supply tech reported at 10:44 PM that D:Q701 was up and running. At 11:10 PM, LRF3 tripped off. It reset. A Pbar Physicist reported at 11:35 PM that he was setting up the lithium lens for stacking. Tuesday April 7 The midnight shift began with the TeV cooling down, with Operations monitoring the stack (14.2E10 antiprotons), with a Pbar Physicist tuning, and with NuMI and MiniBooNE taking beam. Recycler studies ended at 12:25 AM. At 1:07 AM, Operations contacted CHL about the B0 suction pressure rising. A cryo system tech started up some compressors around the TeV ring. Around 1:10 AM, Pbar stopped stacking for about 10 minutes. While operators were investigating, stacking resumed normal operations. At 2:01 AM, Operations reported that the Recycler was left in a strange mode after the study. Operations contacted a Recycler Physicist to find out how to correct the problem so operators could resume transfers. Operators set the system up to take one shot. When the shot went through, a kicker (R:KPS2A) tripped on a “prefire fail to fire” indication. Operations contacted another Recycler Physicist who directed operators to contact an EE Support technician. The tech reported that this was not an issue and to set the Recycler up for a normal transfer. At 2:36 AM, the 750 KeV line Buncher tripped off. It reset. At 3:32 AM, a Linac quadrupole power supply (QPS501) tripped its front panel breaker. It reset. At 3:36 AM, a Booster vacuum pump (B:IPMP02) tripped off and caused MP02 to also trip. Everything reset. At 4:58 AM, Operations received a VESDA alarm for the Booster East tunnel. At 5:08 AM, Operations reported that the Booster RF East Anode Power Supply tripped off. Operators will leave the supply off until the alarm gets resolved. At 5:15 AM, Operations reported that the VESDA levels in Booster East were getting worse. Operators turned off the Booster and prepared it for access by the Fire Department and operators. At 5:19 AM, an EE Support tech reported that a TeV sector F1 beam valve (FIBV7D) would not open. Operations contacted Mechanical Support personnel who power cycled the device and solved the problem. Operations reported at 5:58 AM that Booster was out of access and that operators were turning the machine back on. Operations reported at 6:20 AM a summary of what had happened in the Booster. Two Firemen and an operator accessed the Booster. They found no evidence of a problem. Operations tried to turn on the RF system, but BRF4 tripped off and would not reset and all the other RF stations sparked. Operators bypassed BRF4. Operators then turned on the power supplies; they all turned on except B:V803. Operations received a report from a power supply tech in the Booster West Tower that something in the supply had burned up. Operations determined that BRF4 had caused the VESDA alarm and B:V803 was coincidental. Operations called in two RF technicians to check BRF4 and a power supply tech to check B:V803. At 7:17 AM, the Booster RF West Anode Power Supply (WAPS) tripped off. It reset, but then tripped off again. Operations contacted a Booster RF tech. At 7:30 AM, CDF reported that a UPS battery in the cryo area failed, made a mess, set off alarms, and tripped off the building power. The Fire Department responded and dealt with the mess. The experimenters accessed their hall and began turning on and testing equipment. The Booster power supply tech reported finding a blown disconnect fuse for B:V803 at 7:38 AM. He was still investigating. Operations reported at 8:11 AM that the Recycler had stashed antiprotons twice during the midnight shift. Operations prepared MI for access at 9:31 AM. A power supply technician reported at 10:30 AM that he had replaced the MI-40 Lambda quadrupole power supply. The Neutron Therapy Facility (NTF) reported at 10:36 AM that they had completed treating their last patient for the day. Linac techs reported at 11:44 AM that they were replacing capacitors for Linac Klystron RF stations KRF5, 6, and 7. Booster came out of access at 1:21 PM. The Booster power supply tech reported at 1:33 PM that he was still working on B:V803. Operations reported at 2:08 PM that work on the stacktail equalizer at AP-30 was completed. At 2:35 PM, a Linac quadrupole power supply (L:QPS501) tripped off. It reset. At 2:36 PM, a MI Lambertson (I:LAM10) tripped its wall breaker. A power supply technician was there and investigated. He had the Lambertson back on an hour later. At 2:50 PM, a vacuum tech at TeV sector E34 reported a large amount of water on the floor. Operations reported that the E3SUMP was in high alarm. A Booster power supply tech reported at 3:06 PM that he had brought B:V803 back online. Operators put the Booster into dump mode and a Booster systems specialist established beam to the dump. At 3:15 PM, Operations contacted a power supply tech about problems with three F sector devices: F:M01D. F:MT2WU, and F:MTV3. Operations prepared TeV F sector for access at 3:21 PM. A controls tech reported at 3:43 PM that he had brought the PBKICK front end online. CDF came out of access at 3:56 PM. At 3:59 PM, operators investigating the E3SUMP reported that the pump had no power and the lights were off in the area. The Duty Electrician was contacted. Operations reported at 4:02 PM that the TeV sector E0 crystal collimator was left part way in, but not close enough to interfere with beam. The Operations Proton Source Specialist reported at 4:06 PM that all of the capacitors for the Linac Klystron low-level RF phase loop circuits had been replaced. This successfully eliminated the 130kHz wave from the velocity meter. Operations established beam to MiniBooNE at 5:11 PM. A power supply tech reported at 5:48 PM that F:MT2WU and F:MT3V each had a blown fuse. F:M01D had a regulation problem that resolved itself. He also found that F:MT4W was not regulating well because it had been taken out of voltage mode. He also changed its scale factor. Operations began sending beam to Pbar and NuMI at 7:22 PM. Operations successfully injected protons into the Recycler at 7:38 PM. Operations took the TeV sector A0 kickers out of conditioning mode at 7:48 PM. At 8:57 PM, the 750 KeV buncher tripped off. It reset. At 11:01 PM, BRF4 tripped off for the first of three times in 12 minutes. It reset. At 11:34 PM, the TeV ramp tripped off during the QBS test. Heater Firing Units fired at sectors F1, F2, and F3. The cryo system at F2 went “not ready.” The ramp was back on by 11:51 PM. Operations reported at midnight that the Recycler had stashed antiprotons three times during the evening shift. Wednesday April 8 The midnight shift began with the TeV turn on and cooldown continuing, with Operations monitoring the stash (76.6E10), with Pbar stacking (52.7E10), and with NuMI and MiniBooNE taking beam. At 12:08 AM, CHL reported that the TeV sector F4 cold compressor pump down was not going well. They were contacting the Cryo Coordinator. At 2:41 AM, Operations reported a power glitch. The Main Control Room lights flickered, and the Booster WAPS tripped off along with an Accumulator RF station (ARF1). Everything reset. At 2:44 AM, operators reported a bad smell associated with BRF4. Operations reported at 2:45 AM that CHL had the TeV sector F4 cold compressor on and the Dewar was filling. At 6:08 AM, the Cryo Coordinator reported that there was still a problem with the TeV F4 cryo system. At 7:07 AM, a Pbar Physicist found a core momentum TWT had tripped. It reset. At 7:29 AM, LRF4 tripped off. It reset. NTF reported at 7:30 AM that they were starting to tune up for patient treatment later in the morning. Operations reported at 7:38 AM that the MINOS sump pump was back online. Operations reported at 8 AM that the Recycler had stashed seven times during the midnight shift. Operations began preparing F sector and the Transfer Hall for access at 8:03 AM. Operations loaded an energy conservation time line into the Time Line Generator at 8:07 AM. The Plans for Wednesday The priorities for today are to get the F4 magnets cold and startup the TeV. |
|
More Information Comments and Suggestions |
| last modified 04/10/2009 email Fermilab |
| Security, Privacy, Legal | |